OpenStudy (anonymous):

Solve the equation (2/3)^p=81/16

OpenStudy (anonymous):

ln(81/16)*3/2

OpenStudy (anonymous):

I only get 2.43xxxxxxx

jimthompson5910 (jim_thompson5910):

(2/3)^p = 81/16 (2/3)^p = (3^4)/(2^4) (2/3)^p = (3/2)^4 (2/3)^p = (2/3)^(-4) Since the bases are equal, the exponents are equal. So this means p = -4

OpenStudy (cwrw238):

(2/3)^4 = 16/81 1 / (2/3)^4 = (2/3)^(-4) = 1/16/81 = 81/16 therefore p = -4